# DH-PTAM: A Deep Hybrid Stereo Events-Frames Parallel Tracking And Mapping System

### v1

@@ Abstract
This paper presents a robust approach for a visual parallel tracking and mapping (PTAM) system that excels in challenging environments. Our proposed method combines the strengths of heterogeneous multi-modal visual sensors, including stereo event-based and frame-based sensors, in a unified reference frame through a novel spatio-temporal synchronization of stereo visual frames and stereo event streams. We employ deep learning-based feature extraction and description for estimation to enhance robustness further. We also introduce an end-to-end parallel tracking and mapping optimization layer complemented by a simple loop-closure algorithm for efficient SLAM behavior. Through comprehensive experiments on both small-scale and large-scale real-world sequences of VECtor and TUM-VIE benchmarks, our proposed method (DH-PTAM) demonstrates superior performance compared to state-of-the-art methods in terms of robustness and accuracy in adverse conditions. Our implementation’s research-based Python API is publicly available on GitHub for further research and development: https://github.com/AbanobSoliman/DH-PTAM.
@@
Sensor fusion [1] combines data from multiple sensors to improve a system’s accuracy, reliability, and robustness. It can also reduce computational costs by eliminating the need for redundant sensor data. Different types of sensors can be fused, such as cameras, lidars, radars, and ultrasonics. The algorithm used for fusion can vary, and it typically requires online calibration to ensure accurate and consistent data.
Visual Odometry (VO) is a method that utilizes sensor fusion to estimate the motion of a camera by analyzing the changes in visual features between consecutive frames. Still, it faces challenges, such as difficulties in feature matching when the scene has little texture, the need for a robust feature detector and descriptor, and the problems of relative scale ambiguity and drift [2]. Scale ambiguity refers to the problem of determining the actual scale of the scene. In contrast, drift refers to the accumulation of errors over time that causes the estimated positions to deviate from the true positions. These challenges and limitations must be considered when applying frame-based visual odometry in practical applications [3].
An event camera [4], known as an asynchronous or dynamic vision sensor (DVS), operates on a fundamentally different concept than traditional frame-based cameras. Instead of capturing frames at a constant rate, event cameras output a stream of ”events” that indicate the brightness changes in each pixel. This allows event cameras to operate at high speed, in very low-light conditions, and be more resistant to motion blur [5]. The event-based nature of these cameras also makes them highly suitable for tasks that involve fast-moving objects or scenes with high dynamic range. These characteristics make them an excellent complementary sensor to frame-based visual odometry in adverse conditions such as fast motion, high dynamic range, and low-light environments, where traditional cameras may struggle.
@@
In this paper, we propose a deep hybrid stereo events-frames parallel tracking and mapping system that significantly improves simultaneous localization and mapping accuracy and robustness in dynamic environments. This system combines the advantages of stereo RGB and event cameras, which can capture visual information at high temporal resolution. The use of deep learning techniques in this system allows for the extraction of robust features from the stereo hybrid image and event frames, which improves the accuracy of the feature-matching process and the estimation of the camera pose.
Our main contributions can be summarized as follows:
We propose an end-to-end parallel tracking and mapping (PTAM) approach based on a novel spatio-temporal synchronization of stereo visual frames with event streams (see Fig. 1).
@@
This paper is organized as follows: Section II gives a brief overview of the state-of-the-art SLAM methods. Section III provides an in-detail overview of the proposed method and offers insights of the novel parts of the algorithm. Section IV comprehensively evaluates the algorithm on the most recent VECtor [8] and TUM-VIE [9] benchmarks, along with defining the limitations. Section V summarizes the experiments’ main observations, the proposed method’s behavioral aspects, and the start points for future works.
@@ II Related Work
II-A Conventional visual-SLAM
Simultaneous Localization and Mapping (SLAM) problem has been widely studied in the literature [10], and various techniques have been proposed to solve it. In recent years, learning-based features extraction and description methods [6, 7], and deep learning based approaches [11] have been applied to improve SLAM robustness.
One of the most popular SLAM techniques is the filter-based SLAM using an extended Kalman filter (EKF) [1], or a particle filter [12]. These filters use probabilistic frameworks to estimate the robot’s pose and map. They can handle non-linearities and uncertainties in the system, making them useful for large-scale and highly dynamic environments. Filter-based SLAM has been widely used in applications [13] such as mobile robots, UAVs, and autonomous vehicles.
Another important class of SLAM is graph-based SLAM [14], which uses a factor graph data structure to represent the robot’s poses and the map. Graph-based SLAM requires Sparse Bundle Adjustment (SBA), which uses a non-linear least squares optimization to estimate the robot’s poses and a graph to represent the map. These methods are robust to changes in lighting and viewpoint, making them well-suited for real-world applications. Some popular graph-based SLAM methods include ORB-SLAM [15], Basalt [16], and VINS-Fusion [17].
Loop-closure detection is a fundamental approach to minimize drifts in visual-SLAM, as it allows a system to recognize a previously visited location. Two common approaches to loop-closure detection are mid-level features [18] and bag-of-words [19] representations. Mid-level features are more abstract than low-level features, such as edges and corners, but are not as high-level as object recognition. Deep learning descriptors [20] can be considered mid-level features as they can extract higher-level information from raw data compared to low-level features, such as pixel values, but are not as high-level as features directly related to the task at hand, such as object labels.
II-B Event-aided visual-SLAM
Event-based VO is an emerging form of localization solution that uses event-based cameras to generate measurements of the environment. While the number of sampled frames limits traditional SLAM, event-based SLAM provides high temporal resolution by generating an abundance of measurements, allowing for improved 3D localization and 6D pose estimation. Indirect methods, such as frame-based approaches, extract keypoints from the input data in the front-end. This front-end stage typically involves detecting and matching salient features in the sensory data, such as images or event streams. These keypoints are then passed to the back-end, where state estimation algorithms are used to estimate the robot’s pose and build a consistent map of the environment.
Conversely, direct methods attempt to process all available sensor data, such as individual pixel intensity changes in images (events) or all RGB frame pixels, without any intermediate filtering or feature extraction in the front-end, relying on the back-end to handle the entire data.
Event-aided systems leverage the high-quality representations that events can produce after processing, especially in dynamic and dimmed environments where RGB camera frames fail. Some of the well-known event representations are event image (EI) [21], Time Surfaces (TS) [22], Event Spike Tensor (EST) [23], and recently Event 3-Channel Tensor (E3CT) [24]. Others [4] build the front-end on an Event Generation Model (EGM) [25] or construct motion-compensated event frames (MEF) [26] aided by a gyroscope. Towards a traditional frame reconstruction from events, [27] proposes a Log Intensity Reconstruction (LIR), a model-based method, and [28] proposes Spade-e2vid, a learning-based method.
@@ II-B Event-aided visual-SLAM
| Method | B/A | D/I | EPL | LC | More Information |
@@
| Proposed | A | I | E3CT | ✓ | Stereo (PTAM) + DL |
denotes an Event Pre-processing Layer. denotes Loop-Closure capability. denotes the only method incorporating Deep Learning-aided features.
@@ III-A System Overview
III-B Spatio-temporal Synchronization
Our spatio-temporal synchronization approach (see Fig. 3) considers the general case of global shutter cameras where the exposure time is known. We adopt the constant-time events accumulation window approach in our spatio-temporal events-frames synchronization method.
@@ III-B Spatio-temporal Synchronization
where is the selected stereo keyframe timestamp.
III-C Events-Frames Hybridization Approach
One of the main advantages of our front-end fusion modeling is that it does not rely on any online probabilistic photo-metric matching or alignment approach using filters or cost functions and considers all events polarities . Hence, the computational load of our method lies mainly on the PTAM modules of the optimization-based back-end. The E3CT events pre-processing layer is adopted and modeled as two consecutive filtering kernel convolutions on the event volume of temporal width . The first kernel to filter the time decaying events in the volume, is the -exponential time decay kernel and is modeled as:
@@ III-C Events-Frames Hybridization Approach
Given that the same post-processing operations are applied on the current stereo E3CT frames, the 2D-to-3D-to-2D consecutive inverse-forward projections of the pixels on the E3CT frames to the RGB camera frames can be performed as follows:
where denotes the pixel location in homogeneous coordinates. The term denotes the pixel location alignment correction factor for the RGB and event frames (see Fig. 5) so that the same 3D world point should correspond exactly to the pixel locations . This alignment term is observed to be constant for the same sensor rig with non-varying intrinsic and extrinsic parameters. value can be estimated using an offline optimization process only once on a selected number of frames (the more the accurate) with high confidence feature matches, and this value is given in Section IV for both VECtor and TUM-VIE sequences.
@@
where is the E3CT contribution weight in the current fusion frame. value is dynamic and depends on the scene lighting and texture conditions. It should be set to high values when the RGB camera frame fails to detect features due to adverse conditions and low-textured scenes, and this is the DVS-biased fusion mode. For situations where RGB camera frames can detect reliable scene features with good lighting and enough texture, the value should be low to reduce the amount of extracted features and maintain the back-end processing complexity and latency in reasonable ranges, and this is the APS-biased fusion mode.
Dynamic scenes with challenging and adverse conditions can easily trigger rapid switching between these two fusion modes during long-term navigation. This causes a critical problem during the feature tracking process using conventional low-level feature detectors, such as ORB, SIFT, SURF, BRIEF, and FAST. Accordingly, applying mid-level feature detectors that depend mainly on learning-based architectures could solve this fusion frame modes alternation problem. We employ the learning-based feature extractors and descriptors [6, 7] for their high robustness and feature detection speed. In Fig. 6, we notice the stable tracking of the learning-based features on the hybrid fusion frames in this high dynamic range scenario.
@@ III-D Optimization-based State Estimation
Pose tracking thread. Each map point is projected into the viewing frustum of the anticipated stereo position, and we then look nearby for the match. A valid prediction of the current pose is required for such a projection. By contrasting the descriptions, map points and features are matched. The norm is computed using the binary descriptors of SuperPoint and R2D2. The match is valid if the distance falls below a certain threshold; otherwise, it is ignored. The pose refinement is then applied to recover the current pose knowing the previous one using the following objective function:
@@ IV Evaluation
We perform a thorough, comprehensive evaluation during navigation in real-world, large-scale, and small-scale areas in challenging settings. In subsection IV-A, we compare DH-PTAM with other RGB image-based and event-based/-aided methods on the HDR large-scale sequences of the publicly available dataset VECtor [8] due to its high-quality ground truth values and sensors calibration parameters. In subsection IV-B, we evaluate the small-scale (mocap-) sequences of TUM-VIE [9] to test the quality of the DH-PTAM spatio-temporal synchronization method with degraded event camera calibration parameters. Moreover, the first 45 frames of TUM-VIE sequences suffer a high over-/under-exposure global shutter alternation, which tests the DH-PTAM’s pose estimation stability. We perform a comparative quantitative analysis to evaluate the accuracy of our system in Table II and a qualitative/quantitative analysis in Fig. 7. The accuracy of DH-PTAM is measured with absolute trajectory error (ATE), and relative pose error (RPE) metrics calculated using the baseline SLAM evaluation tool [32].
@@
| Ultimate SLAM (EVIO) [26] | 6.830 | 0.039 | 0.047 | 0.353 | 0.195 | 0.341 | 6.830 | 0.195 | |||||
@@
| DH-PTAM (Stereo Image) | 1.841 | 1.543 | 5.738 | 5.010 | 1.559 | 0.877 | 0.099 | 0.004 | 0.045 | 0.011 | 0.008 | 2.761 | 0.033 |
@@
IMU sensor is included since it is integrated into the front-end and cannot be separated for a fair comparison with EVO, ESVO, and DH-PTAM (ours).
in this ablation case study, the SuperPoint detector is replaced with the R2D2 detector (trained for SLAM tasks), leveraging the GPU performance.
@@ V Conclusion
In this paper, we presented the DH-PTAM system for robust parallel tracking and mapping in dynamic environments using stereo images and event streams. The proposed system builds upon the principles of S-PTAM and extends it with a learning-based approach to handle the sparse and noisy nature of event-based sensors while leveraging the rich information provided by fusion frames. Our experiments demonstrate that DH-PTAM outperforms state-of-the-art visual-inertial SLAM methods, particularly in challenging scenarios such as fast motion, HDR, and occlusions. The proposed system can achieve better performance on a GPU and provides a scalable and accurate solution for 3D reconstruction and pose estimation. Future work includes investigating the potential of integrating inertial navigation sensors, such as IMUs, and exploring the integration of additional deep learning components for improving loop-closure robustness and accuracy. DH-PTAM has the potential to provide robust and accurate 3D mapping and localization, which are crucial for the successful operation of long-term navigation systems.

### v3

@@ Abstract
This paper presents a robust approach for a visual parallel tracking and mapping (PTAM) system that excels in challenging environments. Our proposed method combines the strengths of heterogeneous multi-modal visual sensors, including stereo event-based and frame-based sensors, in a unified reference frame through a novel spatio-temporal synchronization of stereo visual frames and stereo event streams. We employ deep learning-based feature extraction and description for estimation to enhance robustness further. We also introduce an end-to-end parallel tracking and mapping optimization layer complemented by a simple loop-closure algorithm for efficient SLAM behavior. Through comprehensive experiments on both small-scale and large-scale real-world sequences of VECtor and TUM-VIE benchmarks, our proposed method (DH-PTAM) demonstrates superior performance in terms of robustness and accuracy in adverse conditions, especially in large-scale HDR scenarios. Our implementation’s research-based Python API is publicly available on GitHub for further research and development: https://github.com/AbanobSoliman/DH-PTAM.
@@
Simultaneous Localization and Mapping (SLAM) is pivotal in robotics and computer vision, aiming to chart unknown terrains while discerning an agent’s location. Among the key SLAM contributions is the Parallel Tracking and Mapping (PTAM) method [1], which uniquely separates tracking and mapping into parallel threads for augmented efficiency and real-time performance in monocular systems. However, PTAM faced scale ambiguity challenges inherent to monocular SLAM. Addressing this issue, its successor, Stereo Parallel Tracking and Mapping (S-PTAM) [2], leveraged stereo vision to extract depth information directly, eliminating scale ambiguity and fortifying robustness. In recent years, learning-based features extraction and description methods [3, 4], and deep learning based approaches [5] have been applied to improve robustness.
Visual Odometry (VO), an integral component of SLAM, has predominantly depended on conventional cameras to determine motion. However, such methods often fail in high dynamic range (HDR) scenarios where lighting conditions fluctuate [6] (see Fig. 1). Fortunately, the innovation of event cameras, also termed asynchronous or dynamic vision sensors (DVS) [7], offers a groundbreaking solution. Unlike their traditional counterparts that capture frames at fixed intervals, event cameras relay a continuous stream of ”events” showcasing pixel-wise brightness alterations. This not only empowers them to operate at unparalleled speeds and in dimly lit conditions but also markedly diminishes motion blur [8].
With the ability to adeptly handle swift motions, scenes with significant HDR, and challenging lighting, event cameras emerge as superior alternatives, increasing both SLAM and VO’s robustness and precision in environments that would confound traditional cameras. The distinctive event-based design makes these sensors incredibly adapted at tracking fast-moving objects, setting them apart as an invaluable asset to visual odometry, especially under adverse conditions.
@@
This paper proposes a deep hybrid stereo events-frames parallel tracking and mapping system (see Fig. 2) that significantly improves SLAM accuracy and robustness in dynamic environments. This system combines the advantages of stereo RGB and event cameras, which can capture visual information at high temporal resolution. The use of deep learning techniques in this system allows for the extraction of robust features from the stereo hybrid image and event frames, which improves the accuracy of the feature-matching process and the estimation of the camera pose.
The main contributions can be summarized as follows:
We propose an end-to-end parallel tracking and mapping (PTAM) approach based on a novel spatio-temporal synchronization of stereo visual frames with event streams.
@@
This paper is organized as follows: Section II gives a brief overview of the state-of-the-art SLAM methods. Section III provides an in-detail overview of the proposed method and offers insights into the novel parts of the algorithm. Section IV comprehensively evaluates the algorithm on the most recent VECtor [9] and TUM-VIE [10] benchmarks, along with defining the limitations. Section V summarizes the experiments’ main observations, the proposed method’s behavioral aspects, and the start points for future works.
@@ II Related Work
Events-Frames hybridization. Event-aided systems leverage the high-quality representations that events can produce after processing, especially in dynamic and dimmed environments where RGB camera frames fail. Some of the well-known event representations are event image (EI) [12], Time Surfaces (TS) [13], Event Spike Tensor (EST) [14], and recently Event 3-Channel Tensor (E3CT) [11]. Others [7] build the front-end on an Event Generation Model (EGM) [15] or construct motion-compensated event frames (MEF) [16] aided by a gyroscope. Towards a traditional frame reconstruction from events, [17] proposes a Log Intensity Reconstruction (LIR), a model-based method, and [18] proposes Spade-e2vid, a learning-based method.
Indirect methods [19], such as frame-based approaches, extract keypoints from the input data in the front-end. This front-end stage typically involves detecting and matching salient features in the sensory data, such as images or event streams. These keypoints are then passed to the back-end, where state estimation algorithms are used to estimate the robot’s pose and build a consistent map of the environment.
Conversely, direct methods [7] attempt to process all available sensor data, such as individual pixel intensity changes in images (events) or all RGB frame pixels, without any intermediate filtering or feature extraction in the front-end, relying on the back-end to handle the entire data. The proposed method adopts a hybrid approach where all events are directly processed during the events-frames fusion in the front-end, while only the reliable learning-based features from the fusion frames are fed to the back-end (see Fig. 2).
@@
Event-aided visual-SLAM. DH-PTAM builds upon the pioneering work of [7], which introduces a monocular 6-DoF visual odometry model that synergistically integrates events and grayscale frames using a direct approach. In this model, the fusion of events and frames is achieved by applying EGM to the intensity frame, transforming it into a unified reference that aligns with the event stream, termed as brightness increments (event-like). This process is followed by an error minimization cost function, which estimates the extrinsics in real-time for accurate mapping and is augmented by a photometric bundle adjustment for tracking. However, DH-PTAM distinguishes itself from [7] by employing a image-like unified reference in its front-end, harmoniously combining stereo event streams with stereo RGB frames.
Loop-closure detection is paramount in visual-SLAM, effectively minimizing drifts by allowing a system to recognize previously traversed locations. The realm of loop-closure detection predominantly features two methodologies: mid-level features [20] and the bag-of-words model [21]. While mid-level features offer a more nuanced representation than low-level features such as edges and corners, they don’t consider the specificity of high-level features like object recognition. Deep learning descriptors, as referenced in [22], typify mid-level features. They extract more sophisticated information from raw data than low-level features like pixel values, yet they don’t reach the specificity of direct task-related features, for instance, object labels.
| Method | B/A | D/I | EPLa | LCb | More Information |
@@
| Proposed | A | D/I | E3CT | ✓ | Stereo (PTAM) + DLc |
a denotes an Event Pre-processing Layer. b denotes Loop-Closure capability. c denotes the only method incorporating Deep Learning-aided features.
@@ III-A System Overview
III-B Temporal Synchronization Approach
Our temporal synchronization approach (see Fig. 3) considers the general case of global shutter cameras where the exposure time is known. We adopt the constant-time events accumulation window approach where the number of accumulated events during this temporal window is ablated in the qualitative analysis in Fig. 4.
@@ III-B Temporal Synchronization Approach
where the left is the selected stereo keyframe timestamp.
III-C Spatial Hybridization Approach
Leveraging the findings of [25], where Scale-Consistent monocular depth learning is explored, our method is firmly rooted in the depth scale equivalence principle. Rather than just an assumption, this principle is an inference drawn from the inherent properties of closely-spaced APS-DVS sensors. It asserts that the depth scales between frames maintain a consistent relationship. Diverging from the approach in [25], we incorporate the front-end’s depth uncertainty into the back-end’s stereo bundle adjustment process. Here, it acts as an adaptive mechanism, progressively refining depth estimates, thereby boosting system performance. The E3CT events pre-processing layer is adopted and modeled as two consecutive filtering kernel convolutions on the event volume of temporal width (see Fig. 2). The first kernel to filter the time decaying events in the volume, is the -exponential time decay kernel and is modeled as:
@@ III-C Spatial Hybridization Approach
Given that the same post-processing operations are applied on the current stereo E3CT frames (see Fig. 4), the 2D-to-3D-to-2D consecutive inverse-forward projections of the pixels on the E3CT frames to the RGB camera frames can be performed as follows (comparable to Equation (2) in [25]):
where denotes the pixel location in homogeneous coordinates. The term denotes the pixel location alignment correction factor for the RGB and event frames so that the same 3D world point should correspond exactly to the pixel locations . This alignment term is observed to be constant for the same sensor rig with non-varying intrinsic and extrinsic parameters. value can be estimated using an offline optimization process only once on a selected number of frames (the more the accurate) with high confidence feature matches, and this value is given in Section IV for both VECtor and TUM-VIE sequences.
@@
where is the E3CT contribution weight in the current fusion frame. value is calculated online and depends on the scene lighting and texture conditions. It is set to chirp-shaped values when the RGB camera frame fails to detect features due to adverse HDR conditions and low-textured scenes, and this is the DVS-biased fusion mode. For situations where RGB camera frames can detect reliable scene features with good lighting and enough texture, the value is harmonic with the scene lighting conditions according to , and this is the APS-biased fusion mode. To reduce the amount of extracted features and maintain the back-end processing complexity and latency in reasonable ranges, value can be capped at a certain value, as shown in Fig. 6 by setting as an example.
Dynamic scenes with challenging and adverse conditions can easily trigger rapid switching between these two fusion modes during long-term navigation. This causes a critical problem during the feature tracking process using conventional low-level feature detectors, such as ORB, SIFT, SURF, BRIEF, and FAST. Accordingly, applying mid-level feature detectors that depend mainly on learning-based architectures could solve this fusion frame modes alternation problem. We employ the learning-based feature extractors and descriptors [3, 4] for their high robustness and feature detection speed. Fig. 6 elaborates in a real-world case-study, the fusion mode automatic switch using the parameter in response to the rapid intensity fluctuations of the RGB camera frame in HDR scenarios represented by the parameter .
@@ III-D Optimization-based State Estimation
Pose tracking thread. Each map point is projected into the viewing frustum of the anticipated stereo position, and we then look nearby for the match. A valid prediction of the current pose is required for such a projection. By contrasting the descriptions, map points, and features are matched. The norm is computed using the binary descriptors of SuperPoint and R2D2 (see Fig. 7). The match is valid if the distance falls below a certain threshold; otherwise, it is ignored. The pose refinement is then applied to recover the current pose knowing the previous one using the following objective function:
@@ IV Evaluation
We perform a thorough, comprehensive evaluation during navigation in real-world, large-scale, and small-scale areas in challenging settings. In subsection IV-A, we compare DH-PTAM with other RGB image-based and event-based/-aided methods on the HDR large-scale sequences of the publicly available dataset VECtor [9] due to its high-quality ground truth values and sensors calibration parameters. In subsection IV-B, we evaluate the small-scale (mocap-) sequences of TUM-VIE [10] to test the quality of the DH-PTAM spatio-temporal synchronization method with degraded event camera calibration parameters. Moreover, the first 45 frames of TUM-VIE sequences suffer a high over-/under-exposure global shutter alternation, which tests the DH-PTAM’s pose estimation stability. We perform a comparative quantitative analysis to evaluate the accuracy of our system in Table II and a qualitative analysis in Fig. 8. The accuracy of DH-PTAM is measured with absolute trajectory error (ATE), and relative pose error (RPE) metrics calculated using the baseline SLAM evaluation tool [26].
@@
| Ultimate SLAM (EVIO)+ [16] | 6.830 | 0.039 | 0.047 | 0.353 | 0.195 | 0.341 | 6.830 | 0.195 | |||||
@@
| DH-PTAM∗ (Stereo Image) | 1.841 | 1.543 | 5.738 | 5.010 | 1.559 | 0.877 | 0.099 | 0.004 | 0.045 | 0.011 | 0.008 | 2.761 | 0.033 |
@@
+ IMU sensor is included since it is integrated into the front-end and cannot be separated for a fair comparison with EVO, ESVO, and DH-PTAM (ours).
∗ in this ablation case study, the SuperPoint detector is replaced with the R2D2 detector (trained for SLAM tasks), leveraging the GPU performance.
@@ IV-B TUM-VIE small-scale experiments
IV-C Ablation experiments
Front-end depth maps. Qualitatively in Fig. 9, we rigorously assess the front-end of our events-frames fusion technique by comparing its inverse depth maps against established methods such as EVO and EDS. By channeling the stereo fusion frames through the SGM matching method [31], our method capitalizes on direct front-end fusion where the inverse depth maps are derived from events projected onto the RGB frames. Remarkably, the point clouds generated through our method exhibited richer details compared to those of EDS. This affirms the potential of our method to set a new benchmark in event-frame fusion paradigms.
@@ V Conclusion
In this paper, we presented the DH-PTAM system for robust parallel tracking and mapping in dynamic environments using stereo images and event streams. The proposed system builds upon the principles of S-PTAM and extends it with a learning-based approach to handle the sparse and noisy nature of event-based sensors while leveraging the rich information provided by fusion frames. Our experiments demonstrate that DH-PTAM outperforms state-of-the-art visual-SLAM methods, particularly in challenging scenarios such as HDR and occlusions. The proposed system can achieve better performance on a GPU and provides a scalable and accurate solution for 3D reconstruction and pose estimation. Future work includes investigating the potential of improving the spatial synchronization through online optimization and exploring the integration of an adaptive paradigm for choosing the temporal window width to further boost robustness and system latency. DH-PTAM proves to have the potential to provide robust and accurate 3D mapping and localization, which are crucial for the successful operation of long-term navigation systems.